1gram of coprous and oxides was reduced to 0.839g metallic copper. What is the weight of cuprous oxide in sample

Grade:11

1 Answers

Vikas TU
6 years ago
CuO Mass in a given specimen = 40+16 = 56 gm 
Add up to mass of copper in cuprous oxide = 0.839 gm 
Weight of CuO = 0.839/56 = 0.014 gm 
May there be x moles of cu0 and y moles of Cu20 
X * 143 + y * 79.5 = 1 
Weight of CuO = 0.006 * 79.5 = 0.55 gm 
Weight of Cu20 = .45 gm